# Discrete Approximate Circle Bundles

### v2

@@ 4 Effective Computation Of Characteristic Classes
If $\mathcal{U}$ is a finite good open cover and $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\underline{O(2)})$ for some $\varepsilon<\frac{1}{2}$, the outputs $\omega$ and $\tilde{e}$ of the algorithm above are cocycles. If $\varepsilon<\frac{1}{9}$, then $\omega$ and $\tilde{e}$ represent the characteristic classes of $\Pi_{Z}(\Omega)$.
@@
Input: A finite good open cover $\mathcal{U}$ of a closed connected surface $B$ and cocycles $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, $\tilde{e}\in\check{Z}^{1}(\mathcal{U};\mathbb{Z}_{\omega})$.
Output: The twisted Euler number $\tilde{\eta}$ associated to $\tilde{e}$ (up to choice of sign).
  1. 1.
    For $p=2,3$, construct the matrix representation $D_{\omega,p}$ for the twisted boundary operator $\partial_{\omega,p}:\check{C}_{p+1}(\mathcal{U};\mathbb{Z}_{\omega})\to\check{C}_{p}(\mathcal{U};\mathbb{Z}_{\omega})$ with respect to some total orderings on the simplices in each dimension (e.g., lexicographical). Explicitly, the non-zero entries of $D_{\omega,p}$ are given by
    $(D_{\omega,1})_{jkl}^{kl}=\omega_{jk},\quad(D_{\omega,1})_{jkl}^{jl}=-1,\quad(D_{\omega,1})_{jkl}^{jk}=1\\$ (21)
@@
  2. 2.
    Compute the Smith normal form $LD_{\omega,2}R=S$ of $D_{\omega,2}$ over $\mathbb{Z}$ and let $k$ denote the number of zero columns in $S$). If $k=1$, set $\mu_{B,\omega}$ denote the last column of $R$ and proceed to Step 4. Otherwise, let $K$ be the submatrix of $R$ containing the last $k$ columns, and let $B$ equal the last $k$ rows of the product $R^{-1}D_{\omega,3}$.
  3. 3.
    Compute the Smith normal form $L^{\prime}BR^{\prime}=S^{\prime}$ of $B$. Let $v$ denote the last column of $R^{\prime}$, and set $\mu_{B,\omega}=Kv$.
  4. 4.
    Compute $\tilde{\eta}=\sum_{(jkl)\in\mathcal{N}(\mathcal{U})}\tilde{e}_{jkl}\cdot(\mu_{B,\omega})_{jkl}$.
@@
Although Čech theory is typically formulated in terms of cohomology, there is a natural dual notion of Čech homology with the property that $\check{H}_{*}(\mathcal{U};\mathcal{L})\cong H_{*}(B;\mathcal{L})$ for any good open cover $\mathcal{U}$ of a space $B$ and any local system $\mathcal{L}$ of abelian groups (in particular, given a good open cover $\mathcal{U}$, we can define $\check{C}_{k}(\mathcal{U};\mathcal{L})=\text{Hom}_{\mathbb{Z}}(\check{C}^{k}(\mathcal{U};\mathcal{L}),\mathbb{Z})$ and consider the associated homology groups – these groups then coincide with the corresponding singular homology). We can therefore interpret $D_{\omega}$ in the construction above as the matrix representation of a twisted boundary operator whose kernel is generated by the associated twisted fundamental class.
@@
The algorithm above computes the twisted Euler number associated to a twisted Euler class (up to choice of sign).
@@
Since $B$ is a compact connected surface, we have $\text{rank}(H_{2}(B;\mathbb{Z}_{\omega}))=1$, and any generator coincides with the fundamental class up to sign. If $\mathcal{U}$ has no quadruple intersections, then $H_{2}(B;\mathbb{Z}_{\omega})=Z_{2}(B;\mathbb{Z}_{\omega})=\text{ker}\partial_{\omega,2}$, so the last column $r$ of the matrix $R$ above represents a generator of $H_{2}(B;\mathbb{Z}_{\omega})$ (specifically, the vector $r$ corresponds to the simplicial cycle $\sum_{U_{jkl}\in\mathcal{N}(\mathcal{U})^{(2)}}r_{jkl}\cdot U_{jkl}$).
@@ 5 Persistence Of Characteristic Classes
Discrete approximate trivializations constructed from a real-world dataset are unlikely to satisfy the fairly restrictive criteria required for our theoretical guarantees to hold. Due to practical considerations such as non-uniform sampling density or the presence of outliers for which the feature map is ill-defined, we may have reason to believe that some of the local measurements incorporated into our model are more reliable than others. In other cases, we may want to make a deliberate choice to ignore or modify certain parts of the data so that the restricted discrete approximate cocycle representative becomes trivial. Given a family of discrete approximate local trivializations $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ and a witness $\Omega$, we therefore introduce a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ which measures the alignment quality of each pair $f_{j}|_{\pi^{-1}(U_{j}\cap U_{k})}$ and $f_{k}|_{\pi^{-1}(U_{j}\cap U_{k})}$ by the corresponding $\Omega_{jk}$. We then construct an induced simplex-wise filtration $\mathcal{W}$ on $\mathcal{N}(\mathcal{U})$ and compute the persistence of $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ with respect to $\mathcal{W}$. In Section 6, we show how to construct a coordinatization map which is in some sense compatible with the restriction of $\Omega$ to any subcomplex $W^{r}$ in $\mathcal{W}$.
@@ 5.1 Weights Filtration
Suppose $\pi:X\to B$ is a map between metric spaces, $\mathcal{U}$ is an open cover of $B$ and $\Omega\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$. For each $U_{jk}\in\mathcal{N}(\mathcal{U})^{(1)}$, define
$w(U_{jk})=\frac{1}{|\pi^{-1}(U_{jk})|}\sum_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-\Omega_{jk}(\pi(x))f_{k}(x)|$ (22)
Extend $w$ to a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ by setting $w(U_{j})=0$ for all $j$ and inductively defining $w$ for higher simplices by $w(U_{\sigma})=\max_{\sigma^{\prime}\subset\sigma}w(U_{\sigma^{\prime}})$. Order the simplices of $\mathcal{N}(\mathcal{U})$ by (increasing) weight, then by (increasing) dimension, then lexicographically. This ordering induces a simplex-wise filtration $\mathcal{W}^{\Omega}$ of $\mathcal{N}(\mathcal{U})$, called the weights filtration of $\mathcal{N}(\mathcal{U})$ associated with $\Omega$, since all 0-simplices appear first, and any higher simplex appears later than each of its faces.
@@
Suppose $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$. Let $\mathcal{W}^{\Omega}=\{W_{r}\}_{r=1}^{|\mathcal{N}(\mathcal{U})|}$ denote the associated weights filtration. Given a cochain $\Lambda\in\check{C}^{p}(\mathcal{U};\mathcal{F})$ for some presheaf of groups $\mathcal{F}$ and $p\geq 1$, we define the cobirth of $\Lambda$ to be the pair $(r,w(U_{\sigma_{r}}))$, where $r$ is the largest index such that $\Lambda|_{W_{r}}$ is a cocycle. Similarly, we define the codeath of $\Lambda$ to be the pair $(r,w(U_{\sigma_{r}}))$ for $r$ the largest index such that $\Lambda|_{W_{r}}$ is a coboundary.
@@
Suppose $G$ is a PID and $G_{\omega}$ is a local system with stalk $G$. If $\mathcal{U}$ is finite, one can efficiently compute the cobirth and codeath of $\Lambda$ as follows: let $D$ denote the matrix representation of the coboundary operator $\delta_{\omega}:\check{C}^{*}(\mathcal{U};G_{\omega})\to\check{C}^{*}(\mathcal{U};G_{\omega})$ when the simplices are arranged in increasing order according to the weights filtration. Define $\hat{\Lambda}\in\mathbb{R}^{|\mathcal{N}(\mathcal{U})|}$ such that $\hat{\Lambda}_{r}=\Lambda(U_{\sigma_{r}})$ whenever $\sigma_{r}$ is a $p$-simplex and $\hat{\Lambda}_{r}=0$ otherwise. It is easy to see that the cobirth of $\Lambda$ is 1 less than the smallest index $r_{b}$ such that $(D\hat{\Lambda})_{b_{p}}$ is nonzero (if none exists, the cobirth is $|\mathcal{N}(\mathcal{U})|$).
To determine the codeath, write $S=LDR$, where $L$ and $R$ are invertible and $S$ is the Smith normal form of $D$. For $p>0$, let $r_{p}$ denote the largest index of a row of $S$ corresponding to a $p$-simplex and containing a non-zero entry. The codeath of $\Lambda$ is 1 less than the largest index $d_{b}>r_{p}$ such that $(L\hat{\Lambda})_{d_{b}}$ is nonzero (if none exists, then the codeath equals the cobirth).
@@ 5.2 Interpretation
We would like to interpret each filtration stage $W^{r}\in\mathcal{W}$ as the nerve of a modified open cover $\mathcal{U}^{r}$ of a new base space $B^{r}$. For example, removing an edge $(jk)$ from $\mathcal{N}(\mathcal{U})$ effectively separates sets $U_{j}$ and $U_{k}$. For the purposes of coordinatization, there are various choices one could make about how to handle data in $\pi^{-1}(U_{j}\cap U_{k})$, and the appropriate choice depends on context. One option would be to simply remove these points from the dataset – this would be a reasonable choice if the intersection is small and consists mostly of outliers with ill-defined projections. In other cases, one may want to keep the data while modifying $B$ and $\mathcal{U}$. For instance, if the cobirth of a class representative is $\infty$ (or all the weights are fairly small and close together), one may still want to introduce cuts in the base space in order to artificially create a trivial bundle which admits a global coordinate system. With this in mind, we offer a recursive approach for constructing a family $\{B^{r}\}$ of spaces and corresponding open covers $\{\mathcal{U}^{r}\}$ such that $W^{r}=\mathcal{N}(\mathcal{U}^{r})$.
@@
Let $\mathcal{W}^{\Omega}$ denote the weights filtration associated with some $\Omega\in C^{k}(\mathcal{N}(\mathcal{U});\underline{O(2)})$. Suppose $W^{r}\in\mathcal{W}^{\Omega}$ for $r>1$, and we have a space $B^{r}$ and open cover $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j\in J}$ such that $B_{r}=\bigcup_{j\in J}U^{r}_{j}$ and $W^{r}=\mathcal{N}(\mathcal{U}^{r})$. Let $\sigma_{r}\in\mathcal{N}(\mathcal{U})$ denote the simplex which appears in $W^{r}$ but not $W^{r-1}$. Choose some $j\in\sigma_{r}$ (for instance, the last $j$ lexicographically) and define $U^{r-1}_{j}=U_{j}^{r}\cap(\overline{U^{r}_{\sigma_{r}}})^{c}$ and $B^{r-1}=B^{r}-(U^{r}_{j}\cap\partial U^{r}_{\sigma_{r}})$ (for all $k\neq j$, define $U^{r-1}_{k}=U^{r}_{k}$). It is easy to check that $U^{r-1}_{j}\subset B^{r-1}$ is open, $B^{r-1}=\bigcup_{j}U^{r-1}_{j}$ and $W^{r-1}=\mathcal{N}(\mathcal{U}^{r-1})$. In terms of the original space $B$, we have

### v3

@@ 4 Effective Computation Of Characteristic Classes
If $\mathcal{U}$ is a good, finite open cover and $\Omega\in\check{Z}^{1}_{\varepsilon}\left(\mathcal{U};\underline{O(2)}\right)$ for some $\varepsilon<\frac{1}{9}$, the outputs $\omega$ and $\tilde{e}$ of Algorithm 1 represent the characteristic classes associated with $\Pi_{Z}(\Omega)]\in\check{H}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$.
Proof.
By Corollary 4.3, the classes $\det_{*}(\Omega)]$ and $\Delta_{\omega}(\text{abs}(\Omega))$ are respectively the first Stiefel-Whitney class and twisted Euler class of the true circle bundle represented by $\Pi_{Z}(\Omega)$. Expression (27) shows that $\widetilde{e}$ (as defined in 1 above) represents $\Delta_{\omega}(\text{abs}(\Omega))$.
@@
Algorithm 2 (Computation Of Twisted Euler Number).
1:
2:A finite good open cover $\mathcal{U}$ of a closed connected surface $B$ and cocycles $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, $\tilde{e}\in\check{Z}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
3:The twisted Euler number $\tilde{\eta}$ associated to $\tilde{e}$ (up to choice of sign).
4:For $p=2,3$, construct the matrix representation $D_{\omega,p}$ for the twisted boundary operator $\partial_{\omega,p}:\check{C}_{p+1}(\mathcal{U};\mathbb{Z}_{\omega})\to\check{C}_{p}(\mathcal{U};\mathbb{Z}_{\omega})$ with respect to some total orderings on the simplices in each dimension (e.g., lexicographical). Explicitly, the non-zero entries of $D_{\omega,p}$ are given by
$(D_{\omega,1})_{jkl}^{kl}=\omega_{jk},\quad(D_{\omega,1})_{jkl}^{jl}=-1,\quad(D_{\omega,1})_{jkl}^{jk}=1$ (29)
@@
5:Compute the Smith normal form $LD_{\omega,2}R=S$ of $D_{\omega,2}$ over $\mathbb{Z}$ and let $k$ denote the number of zero columns in $S$). If $k=1$, set $\mu_{B,\omega}$ denote the last column of $R$ and proceed to Step 4. Otherwise, let $K$ be the submatrix of $R$ containing the last $k$ columns, and let $B$ equal the last $k$ rows of the product $R^{-1}D_{\omega,3}$.
6:Compute the Smith normal form $L^{\prime}BR^{\prime}=S^{\prime}$ of $B$. Let $v$ denote the last column of $R^{\prime}$, and set $\mu_{B,\omega}=Kv$.
7:Compute $\tilde{\eta}=\sum_{(jkl)\in\mathcal{N}(\mathcal{U})}\tilde{e}_{jkl}\cdot(\mu_{B,\omega})_{jkl}$.
@@
Although Čech theory is typically formulated in terms of cohomology, there is a natural dual notion of Čech homology with the property that $\check{H}_{*}(\mathcal{U};\mathcal{L})\cong H_{*}(B;\mathcal{L})$ for any good open cover $\mathcal{U}$ of a space $B$ and any local system $\mathcal{L}$ of abelian groups (in particular, given a good open cover $\mathcal{U}$, we can define $\check{C}_{k}(\mathcal{U};\mathcal{L})=\text{Hom}_{\mathbb{Z}}(\check{C}^{k}(\mathcal{U};\mathcal{L}),\mathbb{Z})$ and consider the associated homology groups – these groups then coincide with the corresponding singular homology. See [7] for more details). We can therefore interpret $D_{\omega}$ in the construction above as the matrix representation of a twisted boundary operator whose kernel is generated by the associated twisted fundamental class.
@@
If $B$ is a closed connected surface, the algorithm above computes the twisted Euler number associated to a twisted Euler class (up to choice of sign).
@@
Since $B$ is a compact connected surface, we have $\text{rank}(H_{2}(B;\mathbb{Z}_{\omega}))=1$, and any generator coincides with the fundamental class up to sign. If $\mathcal{U}$ has no quadruple intersections, then $H_{2}(B;\mathbb{Z}_{\omega})=Z_{2}(B;\mathbb{Z}_{\omega})=\text{ker}\partial_{\omega,2}$, so the last column $r$ of the matrix $R$ above represents a generator of $H_{2}(B;\mathbb{Z}_{\omega})$ (specifically, the vector $r$ corresponds to the simplicial cycle $\sum_{(jkl)\in\mathcal{N}(\mathcal{U})^{(2)}}r_{jkl}\cdot U_{jkl}$).
@@ 5 Persistence Of Characteristic Classes
Due to practical considerations such as non-uniform sampling density or the presence of outliers for which our feature map is ill-defined, we may have reason to believe that some of the local measurements incorporated into our model are more reliable than others. In other cases, we may want to make a deliberate choice to ignore or modify certain parts of the data so that the restricted discrete approximate cocycle representative becomes trivial. Given a family of discrete approximate local angle functions $\{f_{j}\}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ and a witness $\Omega$, we therefore introduce a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ which measures the alignment quality of each pair $f_{j}|_{\pi^{-1}(U_{j}\cap U_{k})}$ and $f_{k}|_{\pi^{-1}(U_{j}\cap U_{k})}$ by the corresponding $\Omega_{jk}$. We then construct an induced a simplex-wise filtration $\mathcal{W}^{\Omega}$ on $\mathcal{N}(\mathcal{U})$ and compute the persistence of $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ with respect to $\mathcal{W}^{\Omega}$. In Section 6, we show how to construct a coordinatization map which is in some sense compatible with the restriction of $\Omega$ to any subcomplex $W^{r}$ in $\mathcal{W}^{\Omega}$.
@@ 5.1 Weights Filtration
Suppose $\pi:X\to B$ is a function between metric spaces, $\mathcal{U}$ is a good open cover of $B$ and $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})\cong C^{1}(\mathcal{N}(\mathcal{U});O(2))$ witnesses that $\{f_{j}\}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$. For each $(jk)\in\mathcal{N}(\mathcal{U})^{(1)}$, define
$w(jk)=\frac{1}{|\pi^{-1}(U_{jk})|}\sum_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-\Omega_{jk}f_{k}(x)|$ (30)
Extend $w$ to a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ by setting $w(j)=0$ for all $j$ and inductively defining $w$ for higher simplices by $w(\sigma)=\max\limits_{\sigma^{\prime}\subsetneq\sigma}w(\sigma^{\prime})$. Order the simplices of $\mathcal{N}(\mathcal{U})$ by (increasing) weight, then by (increasing) dimension, then lexicographically. This ordering induces a simplex-wise filtration $\mathcal{W}^{\Omega}$ of $\mathcal{N}(\mathcal{U})$, called the weights filtration of $\mathcal{N}(\mathcal{U})$ associated with $\Omega$, since all 0-simplices appear first, and any higher simplex appears later than each of its faces.
@@
Suppose $\Omega\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witnesses that $\{f_{j}\}\in\Gamma_{\alpha}(\pi,\mathcal{U})$. Let $\mathcal{W}^{\Omega}=\{W_{r}\}_{r=1}^{|\mathcal{N}(\mathcal{U})|}$ denote the associated weights filtration. Given a cochain $\Lambda\in\check{C}^{p}(\mathcal{U};\mathcal{F})$ for a presheaf of groups $\mathcal{F}$ and $p\geq 1$, we define the cobirth of $\Lambda$ to be the pair $(r,w(\sigma_{r}))$, where $r$ is the largest index such that $\Lambda|_{W_{r}}$ is a cocycle (assume $p\leq 1$ if $\mathcal{F}$ is a sheaf of non-abelian groups). Similarly, we define the codeath of $\Lambda$ to be the pair $(r,w(\sigma_{r}))$ for $r$ the largest index such that $\Lambda|_{W_{r}}$ is a coboundary.
@@
Suppose $R$ is a PID and $R_{\omega}$ is a local system with stalk $R$. If $\mathcal{U}$ is finite, one can efficiently compute the cobirth and codeath of $\Lambda$ as follows: let $D$ denote the matrix representation of the coboundary operator $\delta_{\omega}:\check{C}^{*}(\mathcal{U};R_{\omega})\to\check{C}^{*}(\mathcal{U};R_{\omega})$ when the simplices are arranged in increasing order according to the weights filtration. Define the vector $\hat{\Lambda}\in\mathbb{R}^{|\mathcal{N}(\mathcal{U})|}$ such that $\hat{\Lambda}_{r}=\Lambda(U_{\sigma_{r}})$ whenever $\sigma_{r}$ is a $p$-simplex and $\hat{\Lambda}_{r}=0$ otherwise. It is easy to see that the cobirth of $\Lambda$ is 1 less than the smallest index $r_{b}$ such that $(D\hat{\Lambda})_{r_{b}}$ is nonzero (if none exists, the cobirth is $|\mathcal{N}(\mathcal{U})|$).
To determine the codeath, write $S=LDR$, where $L$ and $R$ are invertible and $S$ is the Smith normal form of $D$. For $p>0$, let $r_{p}$ denote the largest index of a row of $S$ corresponding to a $p$-simplex and containing a non-zero entry. The codeath of $\Lambda$ is 1 less than the largest index $r_{d}>r_{p}$ such that $(L\hat{\Lambda})_{r_{d}}$ is nonzero (if none exists, then the codeath equals the cobirth).
@@ 5.2 Interpretation
We would like to interpret each filtration stage $W^{r}\in\mathcal{W}$ as the nerve of a modified open cover $\mathcal{U}^{r}$ of a new base space $B^{r}$. For example, removing an edge $(jk)$ from $\mathcal{N}(\mathcal{U})$ effectively separates sets $U_{j}$ and $U_{k}$. For the purposes of coordinatization, there are various choices one could make about how to handle data in $\pi^{-1}(U_{j}\cap U_{k})$, and the appropriate choice depends on context. One option would be to simply remove these points from the dataset – this would be a reasonable choice if the intersection is small and consists mostly of outliers with ill-defined projections. In other cases, one may want to keep the data while modifying $B$ and $\mathcal{U}$. For instance, if the cobirth of a class representative is $\infty$ (or all the weights are fairly small and close together), one may still want to introduce cuts in the base space in order to artificially create a trivial bundle which admits a global coordinate system. With this in mind, we offer a recursive approach for constructing a family of spaces $\{B^{r}\}$ and corresponding open covers $\{\mathcal{U}^{r}\}$ such that $W^{r}=\mathcal{N}(\mathcal{U}^{r})$ for each $r$.
@@
Let $\mathcal{W}^{\Omega}$ denote the weights filtration associated with $\Omega\in\check{C}^{k}(\mathcal{U};\underline{O(2)})$. Suppose $W^{r}\in\mathcal{W}^{\Omega}$ for $r>1$, and we have a space $B^{r}$ and open cover $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j\in J}$ such that $B^{r}=\bigcup_{j\in J}U^{r}_{j}$ and $W^{r}=\mathcal{N}(\mathcal{U}^{r})$. Let $\sigma_{r}\in\mathcal{N}(\mathcal{U})$ denote the simplex which appears in $W^{r}$ but not $W^{r-1}$. Choose some $j\in\sigma_{r}$ (for instance, the last $j$ lexicographically) and define $U^{r-1}_{j}=U_{j}^{r}\cap(\overline{U^{r}_{\sigma_{r}}})^{c}$ and $B^{r-1}=B^{r}-(U^{r}_{j}\cap\partial U^{r}_{\sigma_{r}})$ (for all $k\neq j$, define $U^{r-1}_{k}=U^{r}_{k}$). It is easy to check that $U^{r-1}_{j}\subset B^{r-1}$ is open, $B^{r-1}=\bigcup_{j}U^{r-1}_{j}$ and $W^{r-1}=\mathcal{N}(\mathcal{U}^{r-1})$. In terms of the original space $B$, we have